A projectile is fired in such a way that its horizontal range is equal to its maximum height. What is the angle of projection?


1
Expert's answer
2021-11-07T19:28:42-0500

The range "R" and maximum height "h" are given as follows (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"R = \dfrac{u^2\sin2\theta}{g}\\
h = \dfrac{u^2\sin^2\theta}{2g}"

Equating these expressions, find the angle of projection "\theta":


"\dfrac{u^2\sin2\theta}{g} = \dfrac{u^2\sin^2\theta}{2g}\\
\sin2\theta = \dfrac12\sin^2\theta\\
2\sin\theta\cos\theta = \dfrac12\sin^2\theta\\
4\cos\theta = \sin\theta\\
\tan\theta = 4\\
\theta =\arctan4 \approx 80\degree"

Answer. 80 degrees.